OpenStudy (anonymous):

Solve for x. x + 2(x +1 ) ^1/2 =7

OpenStudy (luigi0210):

Have you made any attempts? Or any idea on how to start?

OpenStudy (anonymous):

I've tried squaring the whole equation to get rid of the square root but it doesn't work.

OpenStudy (luigi0210):

Hmm, did you move everything away from the \((x+1)^{1/2}\)?

OpenStudy (anonymous):

What do you mean? Can you show it?

OpenStudy (luigi0210):

\(\Large x+2(x+1)^{1/2}=7\) into ===> \(\Large (x+1)^{1/2}=\frac{7-x}{2}\) By subtracting x and dividing 2

OpenStudy (anonymous):

Yes, I did try that. What is the next step though?

OpenStudy (luigi0210):

Now you can square everything.. \(\Large ((x+1)^{1/2})^2=\frac{(7-x)^2}{(2)^2}\)

OpenStudy (luigi0210):

Leaving \(\Large x+1=\frac{1}{4}((7-x)^2)\)

OpenStudy (luigi0210):

Multiply everything by 4 \[\Large 4x+4=(49-14x+x^2)\] \[\Large 0=45-18x+x^2\]

OpenStudy (luigi0210):

Now just factor~

OpenStudy (anonymous):

I get it from there now. But can I also ask why didn't it work when I squared it without making (x+1)^1/2 the subject but it worked when I moved everything else to the other side?

OpenStudy (anonymous):

to factorize just break up -18x an such a way that multiplication of coefficients gives 45 and addition gives -18...thaen you can factorize

OpenStudy (luigi0210):

Well I don't know any way of squaring \(\Large x+2(x+1)^{1/2}\).. I'm not sure if that's possible :P

OpenStudy (luigi0210):

You could change the (x+1)^1/2 into a power series but I'm not sure if you know how to do that xD

OpenStudy (anonymous):

Oh okay, then what about solving an equation like this one: \[\sqrt{5 + x} + \sqrt{x} = 5\] Can we still apply the same method?

OpenStudy (anonymous):

write it as.. sqrt5+sqrtx+sqrtx=5 =>2sqrtx=5-sqrt5 now take out squre on both sides...then 4x=25+5-5sqrt5 =>4x=30-5sqrt5 =>x=30/4-5sqrt5/4=15/2-5sqrt5/4

OpenStudy (anonymous):

@fishy13

OpenStudy (anonymous):

But \[\sqrt{a + b} \neq \sqrt{a} + \sqrt{b}\]